OpenStudy (anonymous):

How many hydrogen atoms are in 2.12 mol of propane?

OpenStudy (anonymous):

\[\frac{ 2.12 mol }{ 1 }\frac{ 6.022*10^{23}atoms }{ 1 mole } = 1.28*10^{24}\]